OpenStudy (anonymous):

One number is 5 more than twice another and their sum is 23. The numbers are ______ and _______ . One number is 5 more than twice another and their sum is 23. The numbers are ______ and _______ . @Mathematics

myininaya (myininaya):

n=5+2y n+y=23 (5+2y)+y=23 5+3y=23 3y=23-5 3y=18 y=18/3 y=6 but n=5+2y n=5+2(6) n=5+12 n=17

OpenStudy (hoblos):

6 & 17